OpenStudy (anonymous):

The product of (a − b)(a − b) is a2 − b2. Sometimes Always Never

OpenStudy (anonymous):

never.

OpenStudy (konradzuse):

a^2 -2ab + b^2 so never

OpenStudy (diyadiya):

\[(a-b)(a-b) = (a-b)^2\] and\[(a-b)^2 = a^2 - 2 ab + b^2 \]

OpenStudy (unklerhaukus):

what if a=b=0 or a=b=1

OpenStudy (konradzuse):

to make this simpler you look at the first eq first. So a * a, then you look at the 2nd part so it's b * that same a. then you do the 2nd part. b * a and then b * b.

terenzreignz (terenzreignz):

Or if just b = 0

OpenStudy (konradzuse):

if b = 0 then it would be a^2 if a was 0, then it would be b^2.

OpenStudy (maheshmeghwal9):

so as a summary ur answer is sometimes:)

OpenStudy (konradzuse):

NEVAR!

OpenStudy (maheshmeghwal9):

so i think now u gt:) btw @UnkleRhaukus @terenzreignz gave the best answers

OpenStudy (maheshmeghwal9):

sometimes dude:)

OpenStudy (maheshmeghwal9):

@KonradZuse why do u think like that?

OpenStudy (konradzuse):

Why is it sometimes?

OpenStudy (konradzuse):

The users above said 1 and 0, I already said why 0 doesnt' work.

OpenStudy (maheshmeghwal9):

becoz u r saying never but take b=0 or a=b=1 or a=b=0

OpenStudy (konradzuse):

and if it's 1 then it would either cancel out or it would be a^2 + b^2.

OpenStudy (konradzuse):

0 * anything is 0 so that would make everythign null.

OpenStudy (maheshmeghwal9):

of course cancel out:)

OpenStudy (maheshmeghwal9):

It is a Aptitude question so depends on real domains{inputs}.

OpenStudy (konradzuse):

a^2 - 0 + 0 = a^2 0- 0 + b^2 = b^ 2

OpenStudy (maheshmeghwal9):

so LHS=RHS it is what question wants

OpenStudy (maheshmeghwal9):

u r being confused

OpenStudy (konradzuse):

Possibleh

OpenStudy (maheshmeghwal9):

no need to confuse: - only see this equation is sometimes true for some particular values but without those values it is not possible

terenzreignz (terenzreignz):

@KonradZuse Take a = 4 b = 4 you'll see that the equation holds

OpenStudy (maheshmeghwal9):

yeah that is the good idea.